OpenStudy (anonymous):

root of negative 2x plus 1 − 3 = x Square root of negative 2x plus 1 − 3 + 3 = x + 3 Square root of negative 2x plus 1 = x + 3 Square root of negative 2x plus 1 − 1 = x + 3 − 1 Square root of negative 2 x = x + 2 (Square root of negative 2 x)2 = (x − 4)2 −2x = x2 − 8x + 16 −2x + 2x = x2 + 8x + 16 + 2x 0 = x2 + 10x + 16 0 = (x + 2)(x + 8) x + 2 = 0 x + 8 = 0 x + 2 − 2 = 0 − 2 x + 8 − 8 = 0

OpenStudy (anonymous):

Both solutions are extraneous because they don't satisfy the original equation. What error did Eloise make? She added 2x after squaring both sides. She subtracted 1 before squaring both sides. She factored x2 + 10x + 16 incorrectly. She did not check for extraneous soluti

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

you don't happen to have a screenshot of it?

OpenStudy (anonymous):

hey you go

OpenStudy (jdoe0001):

\(\bf \sqrt{-2x+1}-3=x\\ \sqrt{-2x+1}-3+3=x+3\\ \sqrt{-2x+1}=x+3\\ \color{red}{\sqrt{-2x+1}-1=x+3-1}\\ \sqrt{-2x}=x+2\\ (\sqrt{-2x})^2=(x+2)^2\\ \color{red}{-2x = x^2-8x+16\\ -2x+2x=x^2+8x+16+2x}\\ 0=x^2+10x+16\\ 0=(x+2)(x+8)\) do you notice anything in those lines?

OpenStudy (anonymous):

they add wrong?

OpenStudy (jdoe0001):

hmm actually even more \(\bf \bf \sqrt{-2x+1}-3=x\\ \sqrt{-2x+1}-3+3=x+3\\ \sqrt{-2x+1}=x+3\\ \color{red}{\sqrt{-2x+1}-1=x+3-1}\\ \color{blue}{\sqrt{-2x}=x+2}\\ \color{red}{(\sqrt{-2x})^2=(x-4)^2}\\ \color{blue}{-2x = x^2-8x+16}\\ \color{red}{-2x+2x=x^2+8x+16+2x}\\ 0=x^2+10x+16\\ 0=(x+2)(x+8)\)

OpenStudy (jdoe0001):

\(\bf \sqrt{-2x+1}-1=x+3-1\qquad \textit{can't substract from a radical}\\ \textit{unless the number is inside the radical already}\)

OpenStudy (jdoe0001):

now as far as the others, I gather are typos, notice in the 1st blue line to the red line it goes from x+2 to x-4 for whatever reason notice the 2nd blue to the red line it goes from \(\bf x^2-8x+16 \ \ to \ \ x^2+8x+16\) changing sign on the 8x

OpenStudy (anonymous):

so would it be he factored x2 + 10x + 16 incorrectly.

OpenStudy (jdoe0001):

\(\bf \large {\sqrt{-2x+1}-1\implies \sqrt{-2x+1}-1\\ \quad \\ \sqrt{-2x+1-1}\implies \sqrt{-2x}\\ \quad \\ \sqrt{-2x+1}-1\ne \sqrt{-2x}}\)

OpenStudy (jdoe0001):

hmm if you do FOIL on (x+2)(x+8) it will give you \(\bf x^2+10x+16\)

OpenStudy (jdoe0001):

so that's factored ok

OpenStudy (anonymous):

im so confused would it be She did not check for extraneous solution?

OpenStudy (jdoe0001):

she wanted to isolate the radical and -2x in it, and she wanted to rid of the +1 for that you'd have to square both sides, so you can take out the amount in the radical AND THEN you remove the +1 she used -1 to both sides prematurely, and incorrectly

OpenStudy (anonymous):

o i kind of of get it

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \textit{what she did}&\qquad &\textit{what she should have done}\\ \hline\\ \sqrt{-2x+1}-1=x+3-1&& (\sqrt{-2x+1})^2=(x+3)^2\\ \sqrt{-2x}=x+2&& -2x+1=(x+3)^2\\ &&-2x+1-1=(x+3)^2-1\\ &&-2x = (x+3)^2-1 \end{array}\)